OpenStudy (anonymous):

For problems 13 and 14, assign each letter and a blank space to a number as shown by the alphabet table below. alphabet table 0 = _ 1= A 2 = B 3 = C 4 = D 5 = E 6 = F 7 = G 8 = H 9 = I 10 = J 11 = K 12 = L 13 = M 14 = N 15 = O 16 = P 17 = Q 18 = R 19 = S 20 = T 21 = U 22 = V 23 = W 24 = X 25 = Y 26 = Z use the matrix [1,-2] [-3,7] encode the phrase "ONE QUESTION TO GO"

OpenStudy (anonymous):

@jim_thompson5910 can you help please?

jimthompson5910 (jim_thompson5910):

Sorry for some reason in that last one, I put it as a 9x2 matrix when it *should* be a 2x9 matrix

jimthompson5910 (jim_thompson5910):

ONE_QUESTION_TO_GO turns into the number sequence 15 , 14 , 5 , 0 , 17 , 21 , 5 , 19 , 20, 9 , 15 , 14 , 0 , 20 , 15 , 0 , 7 , 15 which breaks down like this into 2 rows (9 numbers each) 15 , 14 , 5 , 0 , 17 , 21 , 5 , 19 , 20, 9 , 15 , 14 , 0 , 20 , 15 , 0 , 7 , 15

jimthompson5910 (jim_thompson5910):

So you really have to multiply out these 2 matrices \[\large \begin{bmatrix} 1 & -2\\ -3 & 7 \end{bmatrix} \begin{bmatrix} 15 & 14 & 5 & 0 & 17 & 21 & 5 & 19 & 20\\ 9 & 15 & 14 & 0 & 20 & 15 & 0 & 7 & 15 \end{bmatrix} \]

jimthompson5910 (jim_thompson5910):

the same principles of matrix multiplication apply but now you'll have a 2x9 matrix for the answer

OpenStudy (anonymous):

will the numbers be the same but just in a different order? lol bc i just did allll that multiplying for the first go. doing it again would kill -.-

jimthompson5910 (jim_thompson5910):

well in row1, column1 of matrix C you'd multiply row 1 of A with column 1 of B to get 1*15 + (-2)*9 = 15 - 18 = -3 So our 2x9 answer matrix C starts off with 18 blank slots (call them x for now) \[\large C = \begin{bmatrix} x & x & x & x & x & x & x & x & x\\ x & x & x & x & x & x & x & x & x \end{bmatrix} \] and we replace the x in row1, column1 with -3 \[\large C = \begin{bmatrix} -3 & x & x & x & x & x & x & x & x\\ x & x & x & x & x & x & x & x & x \end{bmatrix} \]

jimthompson5910 (jim_thompson5910):

multiply row1 of A with column 2 of B: 1*14 + (-2)*15 = 14 - 30 = -16 this result goes in row1, column2 of C \[\large C = \begin{bmatrix} -3 & -16 & x & x & x & x & x & x & x\\ x & x & x & x & x & x & x & x & x \end{bmatrix} \]

OpenStudy (anonymous):

okay i see. so im gonna try and multiply out this thing and then i will show you for you to check.is that okay?

jimthompson5910 (jim_thompson5910):

sure I can do that

OpenStudy (anonymous):

lol okie dokie sooo i got -3 , -16 , -23 , 0 , -23 , -9 , 5 , 5 , -10 18 , 63 , 83 , 0 , 89 , 42 , -15 , -8 , 45

jimthompson5910 (jim_thompson5910):

I'm getting the same, so nice work

OpenStudy (anonymous):

yay awesome !